0.25 pts Student James Smart heats 5.128 g of sodium carbonate/sodium bicarbonate mixture to a constant mass of 4.084 g, Determine the mass (g) of sodium bicarbonate in the original mixture. Note: employ 3 digits after the decimal point in your answer and do not use units (i.e. g).
Answer #1

2NaHCO3 ---- Na2Co3 + H2o + Co2

The loss in mass id due to the release of H2O and CO2

Loss in mass = 5.128- 4.084 = 1.044g
Molar mass NaHCO3 = 84*2 = 168g
Molar mass CO2 = 44g/mol Molar mass H2O = 18g/mol

168g NaHCO3 will lose 44+18 = 62g in total

Mass loss=1.044g

So,

168/62=x/1.044

Mass of NaHCO3 decomposed =x=2.829

The 5.128g mixture contained 2.829g NaHCO3
Mass of Na2CO3 = 5.128-2.829 =2.299g Na2CO3
% NaHCO3 in mixture =2.829/5.128*100 = 55.167%
%Na2CO3 in mixture = 2.299/5.128*100 = 44.832%